# Duality defect in a deformed transverse-field Ising model

### v3

@@ VII Duality defect in the deformed Floquet TFI model
Now we introduce a duality defect which modifies the deformed TFI Floquet unitary locally around certain sites. An important step is to determine the relative ordering of the non-commuting terms in the Floquet unitary. In particular, given a Trotter choice for the Floquet unitary in the no-defect model, we need a prescription for the ordering of non-commuting terms in the Floquet unitary with the insertion of a duality defect. In [55], this prescription was given in the language of face models of [41]. In order to generalize such prescription to the deformed Floquet TFI model, we first reinterpret this prescription in terms of a half-chain sequential KW transformation, then apply it to obtain the Floquet unitary for the duality defect twisted deformed TFI model. The detailed ordering prescription and derivation for the Floquet unitary can be found in Appendix A.
@@
where $|\beta\rangle$ are the eigenstates of the Floquet unitary $U_{D,\text{dTFI}}$. As some examples, in Figure 18 we plot $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$ for $J=1$, $g=0.3$, $T=4$ with varying deformation strength $\lambda$ and system size $L$. The numerical plateau heights agree well with the values given by Eq. (73). Moreover, the plateau height decreases as the system size increases. Further extrapolation indicates that, as $L\to\infty$, the plateau height approaches zero, indicating its nature of a finite-size effect.
Figure 18: The infinite-temperature auto-correlation function $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$ for $J=1$, $g=0.3$, $T=4$, $L=8,10,12$ and $\lambda=0.1,0.2$. The plateau heights computed using Eq. (73) are indicated by horizontal solid lines.
We now take a closer look at the two stages of $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$: the initial decay stage and the plateau stage. For the parameter choice of $J=1$, $g=0.3$ and $T=4$, numerical analysis suggests that the initial decay of $A_{\infty}^{Y_{i_{0}+1}}$ has a leading $o(\lambda^{3})$ dependence on the deformation strength, as indicated by the data points collapse shown in Figure 19.
Figure 19: The infinite-temperature auto-correlation function $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$ plotted against $\lambda^{3}n$, for $J=1$, $g=0.3$, $T=4$, $L=8,10,12$ and $\lambda=0.1,0.2$.
@@
An important ingredient in the explanations for the existence of a plateau is the nature of eigenstates for the Floquet unitary, which changes drastically as the deformation is turned on. When $\lambda=0$, the Floquet unitary $U_{D,\text{TFI}}$ is quadratic in Majorana fermions within each $\mathbb{Z}_{2}$ charge sector labeled by $\Omega=\pm 1$ (see Eq. (66)). Eigenstates of the Floquet unitary are simple many-particle Fock states $|\Omega,\vec{n}\rangle$, where $n_{i}$ label the occupation numbers for the fermionic modes after a Bogoliubov transformation. Once $\lambda\neq 0$ is turned on, the four-fermion interactions induce hopping processes between different Fock states. Consequently, eigenstates of the Floquet unitary become delocalized in Fock space, namely the eigenstates are now linear combinations of simple Fock states. The degree of this delocalization in Fock space can be quantitatively measured by the inverse participation ratio (IPR), defined as
@@
Figure 20: Plots of IPR and normalized plateau of $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$ as a function of $\lambda$, for $J=1$, $g=0.3$, $T=4$, and $L=10,12$.
Figure 21: Plot of IPR against the normalized plateau of $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$, for $J=1$, $g=0.3$, $T=4$, and $L=10,12$.
@@ Acknowledgements:
Appendix A The ordering convention for Floquet unitaries
@@ Appendix B Decay behavior analysis for the deformed charges at early times
[…]
where
with
and $c_{3}$ is given by
[…]
To show the formal vanishing of $c_{2}$, notice that
The expression for $\left[H^{(1)},Q^{(1)}\right]$ in Eq. (89) is valid given that the 00-th order generator $X^{(0)}$ is well-defined. This is certainly true for the setup of an infinite chain, but raises potential issues for a finite periodic chain. While we do not have an analytical proof for the case of a finite periodic chain, our numerical analysis in Section VI.2 and Section VI.3 does support an $o(\lambda^{3})$ dependence of $A_{\infty}^{Q}(\lambda,t)$ at early times.

### v4

@@ VII Duality defect in the deformed Floquet TFI model
Now we introduce a duality defect which modifies the deformed TFI Floquet unitary locally around certain sites. An important step is to determine the relative ordering of the non-commuting terms in the Floquet unitary. In particular, given a Trotter choice for the Floquet unitary in the no-defect model, we need a prescription for the ordering of non-commuting terms in the Floquet unitary with the insertion of a duality defect. In [55], this prescription was given in the language of face models of [41]. In order to generalize such prescription to the deformed Floquet TFI model, we first reinterpret this prescription in terms of a half-chain sequential KW transformation, then apply it to obtain the Floquet unitary for the duality defect twisted deformed TFI model. The detailed ordering prescription and derivation for the Floquet unitary can be found in Appendix C.
@@
where $|\beta\rangle$ are the eigenstates of the Floquet unitary $U_{D,\text{dTFI}}$. As some examples, in the left panel of Figure 17 we plot $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$ for $J=1$, $g=0.3$, $T=4$ with varying deformation strength $\lambda$ and system size $L$. The numerical plateau heights agree well with the values given by Eq. (62). Moreover, the plateau height decreases as the system size increases. Further extrapolation indicates that, as $L\to\infty$, the plateau height approaches zero, indicating its nature of a finite-size effect.
We now take a closer look at the two stages of $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$: the initial decay stage and the plateau stage. For the parameter choice of $J=1$, $g=0.3$ and $T=4$, numerical analysis suggests that the initial decay of $A_{\infty}^{Y_{i_{0}+1}}$ has a leading $o(\lambda^{3})$ dependence on the deformation strength, as indicated by the data points collapse shown in the right panel of Figure 17.
@@
An important ingredient in the explanations for the existence of a plateau is the nature of eigenstates for the Floquet unitary, which changes drastically as the deformation is turned on. When $\lambda=0$, the Floquet unitary $U_{D,\text{TFI}}$ is quadratic in Majorana fermions within each $\mathbb{Z}_{2}$ charge sector labeled by $\Omega=\pm 1$ (see Eq. (55)). Eigenstates of the Floquet unitary are simple many-particle Fock states $|\Omega,\vec{n}\rangle$, where $n_{i}$ label the occupation numbers for the fermionic modes after a Bogoliubov transformation. Once $\lambda\neq 0$ is turned on, the four-fermion interactions induce hopping processes between different Fock states. Consequently, eigenstates of the Floquet unitary become delocalized in Fock space, namely the eigenstates are now linear combinations of simple Fock states.
Figure 17: Left Panel : The infinite-temperature auto-correlation function $A_{\infty}^{Y_{i_{0}+1}}$ as defined in Eq. (61), plotted against discrete time $n$ for $J=1$, $g=0.3$, $T=4$, $L=8,10,12$ and $\lambda=0.1,0.2$. The plateau heights computed using Eq. (62) are indicated by horizontal solid lines. Right Panel : The infinite-temperature auto-correlation function $A_{\infty}^{Y_{i_{0}+1}}$ plotted against $\lambda^{3}n$ for the same set of parameters.
The degree of this delocalization in Fock space can be quantitatively measured by the inverse participation ratio (IPR), defined as
@@
Figure 18: Plots of IPR in Eq. (63) and normalized plateau of $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$ in Eq. (64) as a function of $\lambda$, for $J=1$, $g=0.3$, $T=4$, and $L=10,12$.
Figure 19: Plot of IPR as defined in Eq. (63) against the normalized plateau of $A_{\infty}^{Y_{i_{0}+1}}(\lambda,n)$ as defined in Eq. (64), for $J=1$, $g=0.3$, $T=4$, and $L=10,12$.
@@ Acknowledgements:
Appendix A Detailed calculations of weak integrability breaking deformations
In this Appendix, we spell out computation details of the weak integrability breaking deformations. We start with the no-defect TFI model discussed in Section VI.2. In the following, we will compute the first-order correction to the Hamiltonian $H^{(1)}=-JQ_{2}^{(1)}$ and see that it reproduces the self-dual deformation in the no-defect TFI model. We will then proceed to compute $Q_{1}^{(1)}$.
Therefore we have
The first-order deformation to the Hamiltonian, $H^{(1)}$, is given by
[…]
where
Substituing this into Eq. (68), we then have
with $h=g/J$ in the convention of Eq. (1).
We now switch to the duality defect twisted TFI model. First of all, as a sanity check, we would like to see that
$\left[Q_{2}^{D,(0)},Q_{1}^{D,(0)}\right]=0~.$ (72)
Direct computation yields
Summing these three terms together gives us
Similarly the generalized current density $J_{22}^{D,(0)}$ is
Next we compute the following anti-commutators
together with
Next, we will compute the deformed charge $Q_{1}^{D,<2}(\lambda)$. From Eq. (31) and Eq. (46), the undeformed duality-twisted charge is given by
Additionally we have
$\begin{split}\text{e}^{\text{i}H(\lambda)t}\mathcal{O}\text{e}^{-\text{i}H(\lambda)t}&=\overline{T}\left\{\text{exp}\left(\text{i}\lambda\int_{0}^{t}dt^{\prime}H^{(1)}_{0}(t^{\prime})\right)\right\}\mathcal{O}_{0}(t)\\ &\times T\left\{\text{exp}\left(-\text{i}\lambda\int_{0}^{t}dt^{\prime}H^{(1)}_{0}(t^{\prime})\right)\right\}\end{split},$ (83)
where
with
and $c_{3}$ is given by
$\begin{split}&-\text{Tr}\left(\int_{0}^{t}dt^{\prime}\left[H^{(1)}_{0}(t^{\prime}),\left(\left[H^{(1)},Q^{(1)}\right]\right)_{0}(t)\right]Q^{(0)}\right)\\ &+\text{i}\text{Tr}\left(\text{e}^{\text{i}H^{(0)}t}\left[H^{(1)},Q^{(1)}\right]\text{e}^{-\text{i}H^{(0)}t}Q^{(1)}\right)\\ =&-\int_{0}^{t}dt^{\prime}\text{Tr}\left(\text{e}^{\text{i}H_{0}(t-t^{\prime})}\left[H^{(1)},Q^{(1)}\right]\text{e}^{-\text{i}H_{0}(t-t^{\prime})}\left[Q^{(0)},H^{(1)}\right]\right)\\ &+\text{i}\text{Tr}\left(\text{e}^{\text{i}H^{(0)}t}\left[H^{(1)},Q^{(1)}\right]\text{e}^{-\text{i}H^{(0)}t}Q^{(1)}\right)\\ =&-\int_{0}^{t}dt^{\prime}\text{Tr}\left(\text{e}^{\text{i}H_{0}t^{\prime}}\left[H^{(1)},Q^{(1)}\right]\text{e}^{-\text{i}H_{0}t^{\prime}}\left[Q^{(0)},H^{(1)}\right]\right)\\ &+\text{i}\text{Tr}\left(\text{e}^{\text{i}H^{(0)}t}\left[H^{(1)},Q^{(1)}\right]\text{e}^{-\text{i}H^{(0)}t}Q^{(1)}\right).\end{split}$ (87)
To show the formal vanishing of $c_{2}$, notice that
The expression for $\left[H^{(1)},Q^{(1)}\right]$ in Eq. (88) is valid given that the 0-th order generator $X^{(0)}$ is well-defined. This is certainly true for the setup of an infinite chain, but raises potential issues for a finite periodic chain. While we do not have an analytical proof for the case of a finite periodic chain, our numerical analysis in Section VI.2 and Section VI.3 does support an $o(\lambda^{3})$ dependence of $A_{\infty}^{Q}(\lambda,t)$ at early times.
Appendix C The ordering convention for Floquet unitaries